(EPS with Convertible Bonds and Preferred Stock) On January 1, 2020, Crocker Company issued 10-year, $2,000,000 face value, 6% bonds, at par. Each $1,000 bond is convertible into 15 shares of Crocker common stock. Crocker's net income in 2020 was $400,000, and its tax rate was 20%. The company had 100,000 shares of common stock outstanding throughout 2020. None of the bonds were converted in 2020.

Question: Compute basic and diluted EPS assuming $400,000 net income, 100,000 weighted avg shares of common stock outstanding and 100,000 shares of convertible 6% $100 par value preferred stock each of which is convertible into 5 shares of common stock
Answer #1

Basic earning per share = 400000/100000 = 4 per share

a) Diluted earning per share = Adjusted net income/Adjusted share = (400000+96000)/(100000+30000) = 3.82 per share

b) Diluted earning per share = 400000/(100000+500000) = 0.67 per share